OpenStudy (anonymous):

A quadratic equation is shown below: x2 + 18x + 76 = 0 Which of the following is the first correct step to write the above equation in the form (x - p)2 = q, where p and q are integers? Add 9 to both sides of the equation Add 5 to both sides of the equation Subtract 5 from both sides of the equation Subtract 9 from both sides of the equation

OpenStudy (anonymous):

please help please im begging

OpenStudy (solomonzelman):

(18/2)²= ? This would be the number you would need to have as the last term.

OpenStudy (anonymous):

so 18?

OpenStudy (solomonzelman):

find that number, and then find the difference between it and between 76.

OpenStudy (solomonzelman):

the last term is NOT 18.

OpenStudy (solomonzelman):

it is (18 ÷ 2) ²

OpenStudy (anonymous):

(18/2)²= 9^2 = 18

OpenStudy (anonymous):

*81

OpenStudy (solomonzelman):

\(\normalsize\color{black}{ (18/2)^2=9^2=9 \times 9 = 81}\) yes....

OpenStudy (solomonzelman):

what do you add to both sides, to make 76 into 81 ?

OpenStudy (anonymous):

so 5

OpenStudy (solomonzelman):

Yes ! +5 to both sides !

OpenStudy (anonymous):

thank you!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (solomonzelman):

then use (a+b)²=a²+2ab+b² (but backwards) you welcome !